# Deep Network Pruning: A Comparative Study on CNNs in Face Recognition

### v1

@@ Abstract
The widespread use of mobile devices for all kind of transactions makes necessary reliable and real-time identity authentication, leading to the adoption of face recognition (FR) via the cameras embedded in such devices. Progress of deep Convolutional Neural Networks (CNNs) has provided substantial advances in FR. Nonetheless, the size of state-of-the-art architectures is unsuitable for mobile deployment, since they often encompass hundreds of megabytes and millions of parameters. We address this by studying methods for deep network compression applied to FR. In particular, we apply network pruning based on Taylor scores, where less important filters are removed iteratively. The method is tested on three networks based on the small SqueezeNet (1.24M parameters) and the popular MobileNetv2 (3.5M) and ResNet50 (23.5M) architectures. These have been selected to showcase the method on CNNs with different complexities and sizes. We observe that a substantial percentage of filters can be removed with minimal performance loss. Also, filters with the highest amount of output channels tend to be removed first, suggesting that high-dimensional spaces within popular CNNs are over-dimensionated.
@@ 1 Introduction
Mobile biometrics can offer secure, user-friendly authentication for many services such as e-commerce, banking, messaging, remote work, social media, education, healthcare, government services, etc. Here, we address face recognition (FR), where Convolutional Neural Networks (CNN), as in many other vision tasks, have become the popular tool [Sundararajan and Woodard, 2018]. Given enough data, CNNs produce classifiers with impressive performance in unconstrained scenarios with high variability. However, the limited computational resources of mobile devices makes that popular CNNs are often unsuitable due to their complexity and large model sizes. This creates the challenge of developing lighter and more efficient models that maintain accuracy without the heavy resource requirements.
@@
In an earlier work [Alonso-Fernandez et al., 2023], we used this method to reduce an already small SqueezeNet network of 1.24M parameters [Iandola et al., 2016; Alonso-Fernandez et al., 2020]. Such work was among the first ones to evaluate network pruning for FR [Caldeira et al., 2024]. This paper extends the study to MobileNetv2 [Sandler et al., 2018] (3.5M parameters) and ResNet50 [He et al., 2016] (23.5M). Such architectures are selected to consider a different amount of model sizes. The purpose and contributions of this work are, therefore, multi-fold. We first summarize literature in network compression, in particular pruning for FR. Then, we study the impact of the employed pruning method on the FR performance of three different-sized CNN architectures. In our experiments, we prune to up to 99% of the filters, maintaining performance until a certain percentage. SqueezeNet and MobileNetv2 allow up to 20-30% pruning, situating them at $\sim$1M parameters. The larger ResNet50 allows up to 99% pruning in some cases which, interestingly, also corresponds to $\sim$1M parameters. We also analyze the effect of pruning in network layers, observing that those with most channels are pruned first.
@@ 2 Related Works
Several light CNNs have been presented across the years [Iandola et al., 2016; Sandler et al., 2018; Zhang et al., 2018; Tan and Le, 2019; Zhang et al., 2020], mainly for generic visual tasks in the context of ImageNet [Russakovsky et al., 2015], using techniques that allow faster processing and fewer parameters such as point-wise convolution, depth-wise separable convolution, variable group convolution, mixed convolution, channel shuffle, and bottleneck layers. Some have been used for FR, either in their original form, or reducing channels or layers for compactness [Chen et al., 2018; Duong et al., 2019; Martinez-Díaz et al., 2019; Yan et al., 2019; Alonso-Fernandez et al., 2020; Boutros et al., 2021]. In doing the latter, however, the number of channels or layers is reduced heuristically just to obtain a more compact network.
@@
Instead of manual adaptation, the work [Boutros et al., 2022] suggested Neural Architecture Search (NAS) to design light FR models named PocketNets. Another increasingly used strategy is network compression. Most biometrics-related compression works have focused on FR, with some others applied to ocular recognition [Almadan and Rattani, 2023], face detection [Jiang et al., 2022], and, to a minor extent, iris, expression, emotion, morphing attacks, or sclera segmentation [Caldeira et al., 2024]. Compression techniques include quantization, knowledge distillation (KD) and pruning. Model quantization converts full-precision (FP) weights and activations (typically in 32 bits) to a lower precision (LP) of 8 bits to up to 1 bit, maintaning the same architecture. LP is less computationally expensive, and several hardware and libraries are optimized for fast 8 bit processing. KD uses a teacher network, usually a complex model that performs well in the proposed task, to guide a ligher student model. The student is trained to produce the output of the teacher via loss functions that minimize the gap between their responses. However, as with quantization, the architecture of KD networks is fixed. In network pruning, on the other hand, redundant parts or connections are eliminated, resulting in a more sparse network. To do so, network weights are usually ordered according to some criteria, and those with the lowest importance are removed until the desired level of sparsity.
@@
Most network compression literature for FR and biometrics in general uses KD [Caldeira et al., 2024]. Apart from our mentioned work [Alonso-Fernandez et al., 2023], few have applied pruning [Polyak and Wolf, 2015; Liu et al., 2022]. The paper [Polyak and Wolf, 2015] proposed two methods to remove low contributing channels: inbound pruning (IP) and reduce and reuse pruning (RRP). IP prunes inbound channels of a layer, wheras RRP prunes outbound channels. The importance of a channel is determined by the variance of its activations. Connections with low variance are deemed as less important, so they can be removed. They also considered hybrid pruning (HP) by first applying RRP followed by IP. The methods were tested on CASIA and LFW datasets, improving runtime up to 2.65 with less than a 2% accuracy drop. The work [Liu et al., 2022] proposed a method to remove channels using a loss function with two terms: a reconstruction error between the feature maps of the baseline network and the pruned one, and a discrimination-aware term that uses the output of intermediate layers as features to do the actual softmax classification. FR experiments used LResNet34E-IR and Mobile-FaceNet as backbones (trained on MS-Celeb-1M) and four well-known benchmark datasets (LFW, CFP-FP, AgeDB-30, and MegaFace), showing negligible performance drops after pruning 25-50% of the channels.
@@ 3.1 Network Pruning
We apply the method for general network pruning of [Molchanov et al., 2019]. It iteratively estimates the importance scores of individual elements based on their effect on the network loss. Then, elements with the lowest scores are pruned, resulting in a more compact network. Given a network with parameters $\textbf{W}=\left\{w_{0},w_{1},...,w_{M}\right\}$ and a training set $\mathcal{D}$ of input $\left(x_{i}\right)$ and output $\left(y_{i}\right)$ pairs $\mathcal{D}=\left\{\left(x_{0},y_{0}\right),\left(x_{1},y_{1}\right),...,\left(x_{K},y_{K}\right)\right\}$, the aim of network training is to minimize the classification error $E$ by solving
@@
Computing $\mathcal{I}_{m}$ for each parameter would demand to evaluate $M$ versions of the network, making the process expensive. This is avoided by approximating $\mathcal{I}_{m}$ in the vicinity of W by its first-order Taylor expansion:
@@
where $g_{m}=\frac{\partial E}{\partial w_{m}}$ are the elements of the gradient $g$. A second-order expansion is also possible via the Hessian of $E$ [Molchanov et al., 2019], but we employ the first-order approximation for faster computation. The gradient $g$ is available from backpropagation, so $\mathcal{I}_{m}^{1}$ can be easily computed. The joint importance of a set of parameters $\textbf{W}_{S}$ (e.g. a filter) can then be obtained as:
@@
The algorithm starts with a trained network, which is pruned iteratively over the same training set. Given a mini-batch, the gradients are computed, the network weights are updated by gradient descent, and the importance of each filter is obtained via Eq. 4. At the end of each epoch, the scores of each filter are averaged over the mini-batches, and those with the smallest importance are removed. The resulting network can be then fine-tuned again to regain potential accuracy losses.
@@ 3.2 Recognition Architectures
We use three CNNs, chosen to account for a different number of parameters or depth (Table 1). They comprise two light architectures (SqueezeNet, MobileNetv2), and a large ResNet50. ResNet networks [He et al., 2016] introduced residual connections, allowing the input from a lower layer to reach a higher layer, bypassing intermediate ones. Followed later by many (including MobileNet), this allows deeper networks and eases training by improving gradient propagation. The authors also introduced bottleneck blocks (Figure 1, left) where high input dimensionality is first reduced with 1$\times$1 point-wise filters, then 3$\times$3 filters are applied over the lower dimensional space, and finally 1$\times$1 filters regain the input dimensionality. The larger 3$\times$3 filters are applied in the reduced dimensionality space, resulting in fewer parameters. We use the ResNet50 variant with 50 convolutional layers, 23.5M parameters, and 83.9 Mb. MobileNetv2 [Sandler et al., 2018] uses depth-wise separable convolutions and inverted residuals to achieve a light architecture (3.5M parameters, 13Mb and 53 convolutional layers). In inverted residuals (Figure 1, center), also called inverted bottleneck blocks with expansion, channel dimensionality is first expanded (instead of reduced) with 1$\times$1 filters. Then, 3$\times$3 filters are applied, but to cope with the increased dimensionality, they are implemented via depth-wise separable convolution for parameter and processing savings. Finally, channel dimensionality is reduced again with 1$\times$1 filters. SqueezeNet [Iandola et al., 2016] is among the smallest architectures proposed within ImageNet (18 convolutional layers, 1.24M parameters, and 4.6 MB), and one of the early attempts to reduce the network parameters and size. It uses squeeze and expand steps, similar to bottlenecks, but with a stack of two layers only called fire module (Figure 1, right). The channel dimensionality is first squeezed with 1$\times$1 point-wise filters, and then expanded with a larger amount of 3$\times$3 and 1$\times$1 filters. SqueezeNet uses late downsampling to keep larger activation maps as much as possible, and it is a sequential CNN without residuals.
Figure 2 shows the architecture of the three networks, from our training environment (Matlab r2023b), with slight changes. The networks use a 113$\times$113 input, achieved by changing the stride of the first convolutional layer from 2 to 1, which keeps the rest of the network unchanged. This allows reusing ImageNet pre-trained networks as starting models for faster convergence [Kornblith et al., 2019]. SqueezeNet uses ReLU activation, and we have added batch normalization (missing in the original implementation), since it helps to improve performance [Alonso-Fernandez et al., 2023]. Bottleneck blocks in ResNet50 include batch normalization and ReLU. MobileNetv2 uses ReLU6 as non-linearity (clipped ReLU), but only after the first two convolutions of inverted bottlenecks.
@@ 3.3 Databases
VGGFace2 [Cao et al., 2018] is used for training and evaluation, comprising 3.31M images of 9131 celebrities (362.6 images per person on average). The images are sourced from the internet, with significant variations in pose, age, ethnicity, lighting, and background. The data is divided into 8631 training classes (3.14M images) and 500 testing classes. A subset of 368 subjects from the test set is defined, VGGFace2-Pose, with 10 images per pose (frontal, three-quarter, profile) and 11040 images. To improve performance, we pre-train the networks with the RetinaFace cleaned subset of MS-Celeb-1M [Guo et al., 2016] (referred to as MS1M), which includes 5.1M images of 93.4K identities. The images are cropped to 113×113 using the five RetinaFace landmarks [Deng et al., 2019b]. While MS1M has more images, its intra-identity variation is limited due to an average of 81 images per person. We address this by pre-training on MS1M, and then fine-tuning on the more diverse VGGFace2. This demonstrated to yield superior performance compared to training on VGGFace2 only [Cao et al., 2018; Alonso-Fernandez et al., 2020]. Some examples of the databases are shown in Figure 3.
@@ 3.4 Training and Verification Protocols
The networks are trained for biometric identification with soft-max and the ImageNet weights as initialization. Training and evaluation follow the VGGFace2 protocol [Cao et al., 2018]. During training, the shortest side of VGGFace2 bounding boxes is resized to 129 pixels and then a random 113$\times$113 crop is taken. MS1M images are directly at 113$\times$113. Additionally, we apply a horizontal random flip. Our optimizer is SGDM with a mini-batch of 128. The learning rate starts at 0.01 and is reduced to 0.005, 0.001, and 0.0001 when the validation loss plateaus. We reserve 2% of images per user in the training set for validation. MS1M users with fewer than 70 images are excluded to avoid fully connected layers dedicated to under-represented classes, resulting in 35016 users and 3.16M images. Our training environment is Matlab r2023b.
We conduct verification experiments with VGGFace2-Pose resizing the shortest image side to 129 pixels, then taking a centre 113×113 crop. VGGFace2-Pose allows same-pose and cross-pose comparisons. Genuine (mated) scores are obtained by comparing each template of a user to his/her remaining ones, avoiding symmetric comparisons. For impostor (non-mated) scores, the first template of a user is used as enrolment and compared with the second template of the next 100 users. Identity templates are created by combining five faces of the same pose [Cao et al., 2018], leading to two templates per user and pose. To test the robustness of the network and the pruning under more challenging conditions, we also use single-image templates. Template vectors are created by averaging face descriptors from the Global Average Pooling layer. To combat pose variation, we average the descriptor of an image with its horizontally flipped counterpart [Duong et al., 2019]. The cosine similarity is used for template comparison. Table 2 shows the total number of scores.
@@
Fig. 7: MobileNetv2 and ResNet50 structure after 26% and 99% pruning respectively. The layer names correspond to the networks’ release of Matlab r2023b. The last layer of MobileNetv2 has 1280 filters (capped for visualization purposes).

### v2

@@ Abstract
The widespread use of mobile devices for all kinds of transactions makes necessary reliable and real-time identity authentication, leading to the adoption of face recognition (FR) via the cameras embedded in such devices. Progress of deep Convolutional Neural Networks (CNNs) has provided substantial advances in FR. Nonetheless, the size of state-of-the-art architectures is unsuitable for mobile deployment, since they often encompass hundreds of megabytes and millions of parameters. We address this by studying methods for deep network compression applied to FR. In particular, we apply network pruning based on Taylor scores, where less important filters are removed iteratively. The method is tested on three networks based on the small SqueezeNet (1.24M parameters) and the popular MobileNetv2 (3.5M) and ResNet50 (23.5M) architectures. These have been selected to showcase the method on CNNs with different complexities and sizes. We observe that a substantial percentage of filters can be removed with minimal performance loss. Also, filters with the highest amount of output channels tend to be removed first, suggesting that high-dimensional spaces within popular CNNs are over-dimensioned.
@@ 1 Introduction
Mobile biometrics can offer secure, user-friendly authentication for many services such as e-commerce, banking, messaging, remote work, social media, education, healthcare, government services, etc. Here, we address face recognition (FR), where Convolutional Neural Networks (CNN), as in many other vision tasks, has become the popular tool [Sundararajan and Woodard, 2018]. Given enough data, CNNs produce classifiers with impressive performance in unconstrained scenarios with high variability. However, the limited computational resources of mobile devices make that popular CNNs are often unsuitable due to their complexity and large model sizes. This creates the challenge of developing lighter and more efficient models that maintain accuracy without heavy resource requirements.
@@
In an earlier work [Alonso-Fernandez et al., 2023b], we used this method to reduce an already small SqueezeNet network of 1.24M parameters [Iandola et al., 2016; Alonso-Fernandez et al., 2020]. Such work was among the first ones to evaluate network pruning for FR [Caldeira et al., 2024]. This paper extends the study to MobileNetv2 [Sandler et al., 2018] (3.5M parameters) and ResNet50 [He et al., 2016] (23.5M). Such architectures are selected to consider different model sizes. The purpose and contributions of this work are, therefore, multi-fold. We first summarize the literature on network compression, in particular pruning for FR. Then, we study the impact of the employed pruning method on the FR performance of three different-sized CNN architectures. In our experiments, we prune up to 99% of the filters, maintaining performance until a certain percentage. SqueezeNet and MobileNetv2 allow up to 20-30% pruning, situating them at $\sim$1M parameters. The larger ResNet50 allows up to 99% pruning in some cases, which, interestingly, also corresponds to $\sim$1M parameters. We also analyze the effect of pruning in network layers, observing that those with most channels are pruned first.
@@ 2 Related Works
Several light CNNs have been presented across the years [Iandola et al., 2016; Sandler et al., 2018; Zhang et al., 2018; Tan and Le, 2019; Zhang et al., 2020], mainly for generic visual tasks in the context of ImageNet [Russakovsky et al., 2015], using techniques that allow faster processing and fewer parameters, such as point-wise convolution, depth-wise separable convolution, variable group convolution, mixed convolution, channel shuffle, and bottleneck layers. Some have been used for FR, either in their original form or reducing channels or layers for compactness [Chen et al., 2018; Duong et al., 2019; Martinez-Díaz et al., 2019; Yan et al., 2019; Alonso-Fernandez et al., 2020; Boutros et al., 2021]. In doing the latter, however, the number of channels or layers is reduced heuristically just to obtain a more compact network. Some other works have specifically addressed optimizations for low-power devices, such as considering mobile face datasets [Cimmino et al., 2022], hierarchical clustering to skip convolutions in image blocks that are similar to each other [Kim et al., 2020], using several shallow CNNs to process each face region separately [Abate et al., 2023], or applying re-lightning and pose correction techniques to fine-tune the recognition network and achieve increased performance in difficult mobile data [Yao et al., 2022].
@@
Instead of manual adaptation, the work [Boutros et al., 2022] suggested Neural Architecture Search (NAS) to design light FR models named PocketNets. Another increasingly used strategy is network compression. Most biometrics-related compression works have focused on FR, with some others applied to ocular recognition [Almadan and Rattani, 2023], face detection [Jiang et al., 2022], and, to a minor extent, iris, expression, emotion, morphing attacks, or sclera segmentation [Caldeira et al., 2024]. Compression techniques include quantization, knowledge distillation (KD) and pruning. Model quantization converts full-precision (FP) weights and activations (typically in 32 bits) to a lower precision (LP) of 8 bits to up to 1 bit, maintaining the same architecture. LP is less computationally expensive, and several hardware and libraries are optimized for fast 8-bit processing. KD uses a teacher network, usually a complex model that performs well in the proposed task, to guide a lighter student model. The student is trained to produce the teacher’s output via loss functions that minimize the gap between their responses. However, as with quantization, the architecture of KD networks is fixed. In network pruning, on the other hand, redundant parts are eliminated, resulting in a more sparse network. To do so, network weights are usually ordered according to some criteria, and those with the lowest importance are removed until the desired level of sparsity.
@@
Most network compression literature for FR and biometrics uses KD [Caldeira et al., 2024]. Apart from our mentioned work [Alonso-Fernandez et al., 2023b], few have applied pruning [Polyak and Wolf, 2015; Liu et al., 2022]. The paper [Polyak and Wolf, 2015] proposed two methods to remove low contributing channels: inbound pruning (IP) and reduce and reuse pruning (RRP). IP prunes inbound channels of a layer, whereas RRP prunes outbound channels. Important channels are determined to be those with a high variance in their activations. Thus, connections with low variance are removed. The methods were tested on CASIA and LFW datasets, improving runtime up to 2.65 with less than a 2% accuracy drop. The work [Liu et al., 2022] proposed to remove channels using a loss function with two terms: a reconstruction error between the feature maps of the baseline network and the pruned one, and a discrimination-aware term that uses the output of intermediate layers as features to do the actual softmax classification. FR experiments used LResNet34E-IR and Mobile-FaceNet as backbones (trained on MS-Celeb-1M) and four well-known benchmark datasets (LFW, CFP-FP, AgeDB-30, and MegaFace), showing negligible performance drops after pruning 25-50% of the channels.
@@ 3.1 Network Pruning
We apply the method for general network pruning of [Molchanov et al., 2019]. It iteratively estimates the importance scores of individual elements based on their effect on the network loss. Then, elements with the lowest scores are pruned. Given a network with parameters $\textbf{W}=\left\{w_{0},w_{1},...,w_{M}\right\}$ and a training set $\mathcal{D}$ of input $\left(x_{i}\right)$ and output $\left(y_{i}\right)$ pairs $\mathcal{D}=\left\{\left(x_{0},y_{0}\right),\left(x_{1},y_{1}\right),...,\left(x_{K},y_{K}\right)\right\}$, the aim of network training is to minimize the classification error $E$ by solving
@@
Computing $\mathcal{I}_{m}$ for each parameter would demand to evaluate $M$ versions of the network. This is avoided by approximating $\mathcal{I}_{m}$ in the vicinity of W by its first-order Taylor expansion:
@@
where $g_{m}=\frac{\partial E}{\partial w_{m}}$ are the elements of the gradient $g$. A second-order expansion is possible via the Hessian of $E$ [Molchanov et al., 2019], but we employ the first-order approximation for faster computation. The gradient $g$ is available from backpropagation, so $\mathcal{I}_{m}^{1}$ can be easily computed. The joint importance of a set of parameters $\textbf{W}_{S}$ (e.g. a filter) can then be obtained as:
@@
The algorithm starts with a trained network, which is pruned iteratively over the same training set. Given a mini-batch, the gradients are computed, the network weights are updated by gradient descent, and the importance of each filter is obtained via Eq. 4. At the end of each epoch, the scores of each filter are averaged over the mini-batches, and those with the smallest importance are removed. The resulting network can then be fine-tuned again to regain potential accuracy losses.
@@ 3.2 Recognition Architectures
We use three CNNs, chosen to account for a different number of parameters or depth (Table 1). They comprise two light architectures (SqueezeNet, MobileNetv2), and a large ResNet50. ResNet networks [He et al., 2016] introduced residual connections, allowing the input from a lower layer to reach a higher layer, bypassing intermediate ones. Followed later by many (including MobileNet), this allows deeper networks and eases training by improving gradient propagation. The authors also introduced bottleneck blocks (Figure 1, left) where high input dimensionality is first reduced with 1$\times$1 point-wise filters, then 3$\times$3 filters are applied over the lower dimensional space, and finally 1$\times$1 filters regain the input dimensionality. The larger 3$\times$3 filters are applied in the reduced dimensionality space, resulting in fewer parameters. We use the ResNet50 variant with 50 convolutional layers, 23.5M parameters, and 83.9 Mb. MobileNetv2 [Sandler et al., 2018] uses inverted residuals to achieve a light architecture (3.5M parameters, 13Mb and 53 convolutional layers). In inverted residuals (Figure 1, centre), also called inverted bottleneck blocks with expansion, channel dimensionality is first expanded (instead of reduced) with 1$\times$1 filters. Then, 3$\times$3 filters are applied, but to cope with the increased dimensionality, they are implemented via depth-wise separable convolution for parameter and processing savings. Finally, channel dimensionality is reduced again with 1$\times$1 filters. SqueezeNet [Iandola et al., 2016] is among the smallest architectures proposed within ImageNet (18 convolutional layers, 1.24M parameters, and 4.6 MB), and one of the early attempts to reduce the network parameters and size. It uses squeeze and expand steps, similar to bottlenecks, but with a stack of two layers, called fire module (Figure 1, right). The channel dimensionality is first squeezed with 1$\times$1 filters, and then expanded with a larger amount of 3$\times$3 and 1$\times$1 filters. SqueezeNet uses late downsampling to keep larger activation maps as much as possible, and it is a sequential CNN without residuals.
Figure 2 shows the architecture of the networks from our training environment (Matlab r2023b), with slight changes. They use a 113$\times$113 input by changing the stride of the first convolutional layer from 2 to 1. This allows keeping the network unchanged and reusing ImageNet pre-trained networks as starting models for faster convergence [Kornblith et al., 2019]. SqueezeNet uses ReLU, and we have added batch normalization (missing in the original implementation), which helps to improve performance [Alonso-Fernandez et al., 2023b]. Bottleneck blocks in ResNet50 include batch normalization and ReLU. MobileNetv2 uses ReLU6 (clipped ReLU), but only after the first two convolutions of inverted bottlenecks.
@@ 3.3 Databases
VGGFace2 [Cao et al., 2018] is used for training and evaluation, having 3.31M images of 9131 celebrities (362.6 images per person on average). The images are sourced from the internet, with significant variations in pose, age, ethnicity, lighting, and background. They are divided into 8631 training classes (3.14M images) and 500 testing classes. A subset of 368 subjects from the test set is defined, VGGFace2-Pose, with 10 images per pose (frontal, three-quarter, profile) and 11040 images. To improve performance, we pre-train the networks with the RetinaFace cleaned subset of MS-Celeb-1M [Guo et al., 2016] (referred to as MS1M), which includes 5.1M images of 93.4K identities. While MS1M has more images, its intra-identity variation is limited due to an average of 81 images per person. We address this by pre-training on MS1M and then fine-tuning on the more diverse VGGFace2, which demonstrated superior performance compared to training on VGGFace2 only [Cao et al., 2018; Alonso-Fernandez et al., 2020]. Some examples of the databases are shown in Figure 3.
@@ 3.4 Training and Verification Protocols
The networks are trained for biometric identification, using ImageNet weights as initialization and soft-max as the activation function in the last layer. Training and evaluation follow the VGGFace2 protocol [Cao et al., 2018]. During training, the shortest side of VGGFace2 bounding boxes is resized to 129 pixels, and then a random 113$\times$113 crop is taken. MS1M images are directly at 113$\times$113. Additionally, we apply a horizontal random flip. We use SGDM with a mini-batch of 128. The learning rate starts at 0.01 and is reduced to 0.005, 0.001, and 0.0001 when the validation loss plateaus. We reserve 2% of images per user in the training set for validation. MS1M users with fewer than 70 images are excluded to avoid fully connected layers dedicated to under-represented classes, resulting in 35016 users and 3.16M images.
We conduct verification experiments with VGGFace2-Pose resizing the shortest image side to 129 pixels, then taking a centre 113×113 crop. Genuine (mated) scores are obtained by comparing each template of a user to his/her remaining ones, avoiding symmetric comparisons. For impostor (non-mated) scores, the first template of a user is compared with the second template of the next 100 users. Identity templates are created by combining five faces of the same pose [Cao et al., 2018], leading to two templates per user and pose. To test the robustness of the network and the pruning under more challenging conditions, we also use single-image templates. Template vectors are created by averaging face descriptors from the Global Average Pooling layer. To combat pose variation, we average the descriptor of an image with its horizontally flipped counterpart [Duong et al., 2019]. The cosine similarity is used for template comparison. Table 2 shows the total number of scores.
@@
Fig. 7: MobileNetv2 and ResNet50 structure after 26% and 99% pruning respectively. The layer names correspond to the networks’ release of Matlab r2023b. The last layer of MobileNetv2 has 1280 filters (capped for visualization purposes). Better if zoomed.